OpenStudy (anonymous):

For which value of instantaneous displacement, the K.E and P.E are equal ??? (x. = amplitude) A. x. /4 B. 3x. /4 C. x. /2 D. x. /undr root 2

OpenStudy (anonymous):

a x/4

OpenStudy (anonymous):

@keshavsarda how?

OpenStudy (anonymous):

this is realeted to shm

OpenStudy (anonymous):

@keshavsarda please explain

OpenStudy (anonymous):

KE=1/2*m*omega^2*v^2 PE=1/2*m*v^2-KE